OpenStudy (idealist10):

Find a second solution y2 for x^2*y"+xy'-y=0; y1=x that isn't a constant multiple of the solution y1. Choose K conveniently to simplify y2.

OpenStudy (idealist10):

So I divided by x^2 and got...\[y''+\frac{ 1 }{ x }y'-\frac{ 1 }{ x^2 }y=0\]

OpenStudy (idealist10):

Now it's in the form of y"+p(x)y'+q(x)y=0.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

I'm not sure if this is the right approach, but notice that you have two "hidden" derivatives: \[\frac{d}{dx}[y']=y''\quad\text{and}\quad\frac{d}{dx}\left[\frac{1}{x}y\right]=\frac{1}{x}y'-\frac{1}{x^2}y\] so the ODE is equivalent to \[\frac{d}{dx}\left[y'+\frac{1}{x}y\right]=0\] Integrating both sides yields \[y'+\frac{1}{x}y=C_1\] which is linear in \(y\). If you were to look for an integrating factor, you'd find that \(\mu(x)=x\): \[\begin{align*}xy'+y&=C_1x\\\\ \frac{d}{dx}[xy]&=C_1x\\\\ xy&=\frac{1}{2}C_1x^2+C_2\\\\ y&=\frac{1}{2}C_1\underbrace{x}_{y_1}+C_2\underbrace{\frac{1}{x}}_{y_2} \end{align*}\]

OpenStudy (anonymous):

Another method (again, not sure if it's what you're looking for): the given ODE is of the Euler-Cauchy form, \[x^2y''+xy'-y=0\] Setting \(y=x^r\), you have the derivatives \(y'=rx^{r-1}\) and \(y''=r(r-1)x^{r-2}\). Subbing these in, you get \[r(r-1)x^r+rx^r-x^r=0~~\implies~~r(r-1)+r-1=(r+1)(r-1)=0\] which has roots \(r=\pm1\). Here you find that the fundamental solutions would agree with the ones found before, \(\left\{x,\dfrac{1}{x}\right\}\).

OpenStudy (anonymous):

Another method, this time using reduction of order (probably what you wanted in the first place, now that I think about it). Assume you have a solution of the form \(y=y_1v=xv\), then \(y'=xv'+v\) and \(y''=xv''+2v'\). Subbing into the ODE, you have \[\begin{align*}x^2(xv''+2v')+x(xv'+v)-xv&=0\\ x^3v''+3x^2v'&=0\\ xv''+3v'&=0\end{align*}\] Substituting \(t=v'\), you get the linear ODE, \[xt'+3t=0\]

OpenStudy (idealist10):

I love your first method of solving this problem. Thank you so much!